A 400-V battery pack is used to accelerate an electric car over a time interval of 9.5 s. If 2427 C of charge flow through the battery pack during this acceleration period, determine the horsepower of the car's engine. Please show work if possible.Please answer in hp, not kw.
Answer #1

voltage v = 400v

charge q = 2427 c

time t = 9.5 sec

power p =w/t = vq/t =400x2427/9.5 = 102189.5watts= 137 hp

1 horse power = 746 w   =102189.5/746 = 137 hp